OpenStudy (anonymous):

Which of these is a literal equation? ten minus six equals four three squared plus four equals ten plus three x squared minus seven equals y plus four three time eight minus ten equals three time negative two

OpenStudy (goldphenoix):

To answer this question, you want to know what a literal equation is. Do you know what a literal equation is?

OpenStudy (anonymous):

nope. -.-

OpenStudy (goldphenoix):

Alright, a literal equation is an equation with more than one variable. Do you know what a variable is?

OpenStudy (goldphenoix):

Yes, no, maybe?

OpenStudy (goldphenoix):

Anyway, it would be easier to answer this question if you write each of the multiple choices in standard notation.

OpenStudy (goldphenoix):

ten minus six equals four = 10 - 6 = 4 three squared plus four equals ten plus three = 3^2 + 4 = 10 + 3 x squared minus seven equals y plus four = x^2 - 7 = y + 4 three time eight minus ten equals three time negative two = 3 * 8 - 10 = 3 * -2

OpenStudy (anonymous):

so whats the answer im in a hurry

OpenStudy (goldphenoix):

Do you know what a variable is? If you know what a variable is, then you already found your answer.

OpenStudy (goldphenoix):

Since your in a hurry, I'm just going to give you a review on what a variable is. Variable - A symbol for a number we don't know yet. It is usually a letter like x or y. Example: in x + 2 = 6, x is the variable

OpenStudy (goldphenoix):

Since a literal equation has more than 1 variable, then the answer is an equation that has 1 or more varaible.

OpenStudy (goldphenoix):

Did you find your answer?

OpenStudy (anonymous):

no

OpenStudy (goldphenoix):

ten minus six equals four = 10 - 6 = 4 three squared plus four equals ten plus three = 3^2 + 4 = 10 + 3 x squared minus seven equals y plus four = x^2 - 7 = y + 4 three time eight minus ten equals three time negative two = 3 * 8 - 10 = 3 * -2 Which answer choices has more than 1 variable?

OpenStudy (anonymous):

2 I believe

OpenStudy (goldphenoix):

Your saying that "three squared plus four equals ten plus three = 3^2 + 4 = 10 + 3" has more than 1 variable?

OpenStudy (goldphenoix):

This is the only hint that I can give you. An answer choice that has more than 1 variable. Look at each answer choices carefully.

OpenStudy (anonymous):

ummm

OpenStudy (goldphenoix):

Hmm? You still don't get it?

OpenStudy (anonymous):

nope

OpenStudy (goldphenoix):

How many variable(s) are there in 10 - 6 = 4 ?

OpenStudy (anonymous):

1

OpenStudy (goldphenoix):

Where's the variable in that equation?

OpenStudy (anonymous):

6

OpenStudy (goldphenoix):

Well. As I said 30 minutes ago, Variable - A symbol for a number we don't know yet. It is usually a letter like x or y. Example: in x + 2 = 6, x is the variable

OpenStudy (goldphenoix):

6 is not a variable. 6 is a constant. A variable is a letter, not a number.

OpenStudy (goldphenoix):

So how many variable(s) are there in 10 - 6 = 4?

OpenStudy (naga447):

It is the 3rd one down...x squared minus seven equals y plus four = x^2 - 7 = y + 4